### v1

@@ IV.2 Part of a Trend?
We argue that the clear-cut event analyzed in the present letter is not a one-off, but rather part of a trend, where Ref. [67] presents relevant evidence from the dayside polar ionosphere. Indeed, the turbulent Hall channels that may frequent the vicinity of diffuse aurorae in general are highly localized in time and space, like most research into the “spiky” nature of electric field enhancements can attest [54]. To support such a general connection between magnetospheric wave-particle energy and subsequent modulations to the icebear echo detection rates, we have fallen back on inferences made from statistics.
@@
In Figure 6a–b) we show the resulting median values for the echo detection rates and wave power respectively, in 15 geomagnetic activity bins. In Panel a) we observe a distinct kink in the curve at an SME-index value of around 150 nT, below which the echo detection rate is flat (indicating a contamination by meteor trail echoes [68]). For bins above this kink, the bin-median echo detection rate is perfectly correlated ($\rho=1$) with the SME-index. Crucially, the linear slope ($1.71\pm 0.11$ in a log-log scale) is roughly consistent with the slope exhibited by the wave power data (Figure 6b), $2.23\pm 0.31$.
The two quantities exhibit a similar response to enhancements in the SME-index. This facilitates the 1:1 relationship shown in Figure 6c), where we display the same geomagnetic activity bins in a scatterplot (excluding the gray datapoints in Panel a), and a log-log linear fit with slope 1 is shown with a red, dashed line. Lastly, with a green-blue circle, we show the median values derived from the 12 May 2021 event, with errorbars denoting upper/lower quartile distributions.
Figure 6a-b) show two widely different geophysical quantities – one in the ionosphere and one in the magnetosphere – that nevertheless respond in a similar fashion to a third independent variable, the SME-index. Indeed, that index is itself an unambiguous measurement from the ground of the ionosphere’s high-latitude Hall currents (the auroral electrojets) [69, 70]. Enhancements in Hall conductance are, in turn, driven by energetic particle precipitation [71, 72], by virtue of providing ionization and driving electric field enhancements. As we detail above (and illustrate in Figure 5), the same two quantities ultimately drive Farley-Buneman turbulence.
In empirical terms, the diffuse and pulsating aurorae constitute the majority of the total energy input into the nightside ionosphere [73], and so enhancements in the Hall currents are largely driven by diffuse and pulsating aurorae [22, 74]. There, ECH waves constitute a vital part of the acceleration mechanisms [50, 11, 17, 44].
@@
The above chain of argument explains why the SME-index is, on average, perfectly correlated with icebear’s echo count. The SME-index is a global quantity, and so the echo count is not expected to track it during individual events. The average number of echoes per time-interval will, given a large enough sample, be perfectly correlated with the SME-index.
@@ Acknowledgements
This work is supported in part by Research Council of Norway (RCN) grant [324859]. We acknowledge the support of the Canadian Space Agency (CSA) [20SUGOICEB], the Canada Foundation for Innovation (CFI) John R. Evans Leaders Fund [32117], the Natural Science and Engineering Research Council (NSERC), the Discovery grants program [RGPIN-2019-19135], the Digital Research Alliance of Canada [RRG-4802], and basic research funding from Korea Astronomy and Space Science Institute [KASI2024185002]. Science data of the ERG (Arase) satellite were obtained from the ERG Science Center operated by ISAS/JAXA and ISEE/Nagoya University (https://ergsc.isee.nagoya-u.ac.jp/index.shtml.en. This includes Lv.3 MEP-e (DOI 10.34515/DATA.ERG-02003), Lv.2 PWE/OFA (DOI 10.34515/DATA.ERG-08000), Lv.2 PWE/HFA (DOI 10.34515/DATA.ERG-10000), and Lv.2 MGF (DOI 10.34515/DATA.ERG-06001). icebear 3D echo data for 2020, 2021 is published with DOI 10.5281/zenodo.7509022. Supermag data can be accessed at https://supermag.jhuapl.edu/mag/. The SymH-index from NASA’s omni service can be accessed at https://omniweb.gsfc.nasa.gov/. trex optical data can be accessed at https://data.phys.ucalgary.ca/ [78]. MFI is thankful to F Lind, Y Jin, W Miloch, and L Clausen for stimulating discussions.

### v2

@@ IV.2 Part of a Trend?
We argue that the clear-cut event analyzed in the present letter is part of a trend, where Ref. [66] presents relevant evidence for this trend from the dayside polar ionosphere. Indeed, the turbulent Hall channels that may frequent the vicinity of diffuse aurorae in general are highly localized in time and space, like most research into the “spiky” nature of electric field enhancements can attest to [54]. To support such a general connection between magnetospheric wave-particle energy and subsequent modulations to the icebear echo detection rates, we have fallen back on inferences made from statistics.
@@
In Figure 6a–b) we show the resulting median values for the echo detection rates and wave power respectively, in 15 geomagnetic activity bins. In Panel a) we observe a distinct kink in the curve at an SME-index value of around 150 nT, below which the echo detection rate is flat (indicating a contamination by meteor trail echoes [67]). For bins above this kink, the bin-median echo detection rate is perfectly correlated ($\rho=1$) with the SME-index. The linear slope ($1.71\pm 0.11$ in a log-log scale) is roughly consistent with the slope exhibited by the wave power data (Figure 6b), $2.23\pm 0.31$.
The two quantities exhibit a similar response to enhancements in the SME-index. This facilitates the 1:1 relationship shown in Figure 6c), where we display the same geomagnetic activity bins in a scatterplot (excluding the gray datapoints in Panel a as well as the equivalent bins in panel b), and a log-log linear fit with slope 1 is shown with a red, dashed line. Lastly, with a green-blue circle, we show the median values derived from the 12 May 2021 event, with errorbars denoting upper/lower quartile distributions.
Figure 6a-b) show two widely different geophysical quantities – one in the ionosphere and one in the magnetosphere – that nevertheless respond in a similar fashion to a third independent variable, the SME-index. Indeed, that index is itself an unambiguous measurement from the ground of the ionosphere’s high-latitude Hall currents (the auroral electrojets) [68, 69]. Enhancements in Hall conductance are, in turn, driven by energetic particle precipitation [70, 71], by virtue of providing ionization and causing electric field enhancements. As we detail in the present letter (and illustrate in Figure 5), the same two quantities are able to drive Farley-Buneman turbulence.
In empirical terms, the diffuse and pulsating aurorae constitute the majority of the total energy input into the nightside ionosphere [72], and so enhancements in the Hall currents there are largely driven by diffuse and pulsating aurorae [22, 73].
@@
The above chain of argument explains why $\rho=1$ in Figure 6a): Global though it may be, and thus not expected to correlate with the echo rates during individual events, the SME-index encapsulates the average response of the icebear data in the face of an externally driven magnetosphere.
@@ Acknowledgements
This work is supported in part by Research Council of Norway (RCN) grant [324859]. We acknowledge the support of the Canadian Space Agency (CSA) [20SUGOICEB], the Canada Foundation for Innovation (CFI) John R. Evans Leaders Fund [32117], the Natural Science and Engineering Research Council (NSERC), the Discovery grants program [RGPIN-2019-19135], the Digital Research Alliance of Canada [RRG-4802], and basic research funding from Korea Astronomy and Space Science Institute [KASI2024185002]. Science data of the ERG (Arase) satellite were obtained from the ERG Science Center operated by ISAS/JAXA and ISEE/Nagoya University (https://ergsc.isee.nagoya-u.ac.jp/index.shtml.en. This includes Lv.3 MEP-e (DOI 10.34515/DATA.ERG-02003), Lv.2 PWE/OFA (DOI 10.34515/DATA.ERG-08000), Lv.2 PWE/HFA (DOI 10.34515/DATA.ERG-10000), and Lv.2 MGF (DOI 10.34515/DATA.ERG-06001). icebear 3D echo data for 2020, 2021 is published with DOI 10.5281/zenodo.7509022. Supermag data can be accessed at https://supermag.jhuapl.edu/mag/. The SymH-index from NASA’s omni service can be accessed at https://omniweb.gsfc.nasa.gov/. trex optical data can be accessed at https://doi.org/10.11575/4P8E-1K65. MFI is thankful to F Lind, Y Jin, W Miloch, and L Clausen for stimulating discussions.
See Supplemental Material [url] for additional frequency-dependent correlation akin to Figure 3, as well as a brief discussion of lower-band chorus waves in the same context. The Supplemental Materials includes Refs. [77, 78, 79, 80, 81].